OpenStudy (anonymous):

The age of the Earth is approximately 1.3 1017 seconds, and one year is approximately 3.2 107 seconds. Which of the following is closest to the age of the Earth in years? A. 2.5 109 B. 4.1 109 C. 1.9 1010 D. 2.5 1011 E. 4.1 1011

OpenStudy (anonymous):

pls reply me, i m waiting

OpenStudy (anonymous):

\[\frac{1.3*10^{17}}{3.2*10^7}= \frac{1.3}{3.2} *\frac{10^{17}}{10^7}\]